### v1

# A simplified directional KeRF algorithm

Abstract.
Random forest methods belong to the class of non-parametric machine learning algorithms. They were first introduced in 2001 by Breiman and they perform with accuracy in high dimensional settings. In this article, we consider, a simplified kernel-based random forest algorithm called simplified directional KeRF (Kernel Random Forest). We establish the asymptotic equivalence between simplified directional KeRF and centered KeRF, with additional numerical experiments supporting our theoretical results.
The first author was funded by the PhD scholarship PON-DOT1303154-3, the second author
1. Introduction
Random forest algorithms are ensemble machine learning methods for regression and classification tasks. The class of random forest algorithms performs with high accuracy in high-dimensional data avoiding in general overfitting. It has become one of the most important practical data analysis tools in various active research fields, mainly because of its high performance even in sparse data sets.
Decision trees and random forests have been used in many active research fields in modern sciences and belong to the family of non-parametric machine learning algorithms. In particular, they have been used in economics [Yoo21], biology [BJKK12], linguistics [Gri19], bio-informatics [DKN15], and genetics [CI12] to name a few. In practice random forest-type algorithms with many variations are commonly used, since usually, few parameters need to be tuned for training the model.
The random forest algorithm was originally introduced by Breiman in a seminar paper in [BFOS84]. Following the previous work of Ho [Ho95], where it was proposed a method following stochastic modeling for creating multiple tree classifiers, and Amit and Geman [AG97], who introduced a new approach to shape recognition based on tree classifiers as well, Breiman introduced the CART (Classification and Regression Trees) method. Random forests are widely used in many various research domains since only a few parameters need to be tuned [LW02], [GPT08] but still their theoretical properties are under active research [LJ06], [BD10]. For a review of the theoretical and practical aspects of the theory up to 2016, the authors refer to [BS16].
An important distinction within the class of random forest algorithms lies in the construction of each individual tree. The original random forest algorithm proposed by Breiman [BFOS84] uses the data set to define the estimator through the CART method. Significant research also addresses the consistency of data-dependent algorithms [MH14], [Wag14], [SBV15]. On the contrary, another class of algorithms are those designed independently of the data set the so-called purely random forests [BDL08], [Sco16a], [DMF13]. In 2012, Biau examined thoroughly a specific model proposed by Breiman belonging to the class of purely random forests, the centered random forest [Bia12]. In this in-depth analysis, Biau proved the consistency of the algorithm and the dependence of the convergence rate only on the strong variables of the feature space and not on the noisy ones.
Of particular interest to researchers is the rate of convergence of the various random forest algorithms. In particular, in [Bia12], Biau provided an upper bound on the rate of the consistency of the centered random forest over the class of the Lipschitz functions of $\mathcal{O}\left(n^{-\frac{1}{d\frac{4}{3}\log{2}+1}}\right),$ where $n$ is the size of the sample data and $d$ the dimension of the feature space. Moreover, Klusowski in [Klu18] provided an improved rate of convergence to $\mathcal{O}\left((n\log^{\frac{d-1}{2}}{n})^{-(\frac{1+\delta}{d\log{2}+1})}\right),$ where $\delta$ is a positive constant that depends on the dimension of the feature space $d$ and converges to zero as $d$ approaches infinity. The aforementioned rate of convergence is sharp although it fails to reach the minimax rate of convergence in the class of the Lipschitz functions [YB99] $\mathcal{O}\left(n^{\frac{-2}{d+2}}\right).$
An important direction in manipulating machine learning algorithms is through kernel theory. Already Breiman in [Bre00] introduced the connection between kernel methods and random forests where he proved the equivalence between the estimator of a purely random forest and a kernel-type algorithm. In 2006 Geurts et al. in [GEW06] formalized the idea by defining the extremely randomized trees method that has a kernel characterization.
In 2016 Scornet introduced the Kernel Random Forest method (KeRF) [Sco16b]. By slightly modifying the definition of random forests, Scornet suggested a new estimator that has a kernel representation. The new KeRF estimators are not of the form of the well-studied Nadaray-Watson kernels with bandwidth and a more careful analysis is necessary. In [Sco16b], the expressions for the centered and uniform KeRF algorithms were explicitly defined, along with upper bounds on their rate of convergence. Additionally, due to the kernel representation of the algorithm, the exact formula of the infinite-centered KeRF can be computed (for a more detailed description, see 2.1).
In this article, we introduce a variation of the centered random forest algorithm called the simplified directional algorithm. Our objective is to define a partition of the feature space independently of the data set by simplifying the centered method. We derive the corresponding kernel representation of the algorithm and we prove that, asymptotically, as the number of trees tends to infinity, the centered KeRF and the simplified directional KeRF coincide. Moreover, we support our theoretical results with various experiments, comparing the $L_{2}$-error and the variance of the centered KeRF and the simplified directional algorithm for different numbers of trees.
In the first part of the article, we provide the notation and the definitions of random forests and Kernel Random Forests (KeRF). We introduce the simplified directional algorithm and prove that the probability of two points being connected in a tree of depth $k$ is the same for both KeRF algorithms. In the second part of the paper, we examine numerical experiments on various data sets for different target functions.
2. Notation
A fundamental problem in machine learning is to learn the relationship between input features ($\mathbf{X}$) and an output target variable ($Y$) from a set of training examples. Specifically, given a training sample $(\mathbf{X}_{1},Y_{1}),(\mathbf{X}_{2},Y_{2}),...,(\mathbf{X}_{n},Y_{n}),$ of independent random variables, where each $\mathbf{X}_{i}$ is a $d$-dimensional vector in $[0,1]^{d}$ and $Y_{i}$ is a real-valued output, the goal is to estimate the function $m(\mathbf{x})=\mathbb{E}(Y|\mathbf{X}=\mathbf{x})$. In other words, we want to predict the expected value of $Y$ for any new input vector $\mathbf{x}$, based on the training data. In classification problems, $Y$ takes on values from a finite set of class labels.
We typically assume that the training sample $(\mathbf{X}_{i},Y_{i})$ are independent and identically distributed random variables drawn from the same unknown joint distribution $\mathbb{P}(\mathbf{X},Y)$.
The learning algorithm must then use just these $n$ samples to construct a function $m_{n}:[0,1]^{d}\rightarrow\mathbb{R}$ that is a good approximation to the conditional expectation $m(\mathbf{x})=\mathbb{E}(Y|\mathbf{X}=\mathbf{x})$ for $x\in[0,1]^{d}.$
Next, we consider the basic framework of the random forest algorithm, which is an ensemble learning method that combines multiple decision trees to improve predictive performance and reduce overfitting.
A random forest is a finite collection (average) of $M-$ independent, regression random trees designed with respect to the identical distributed random variables $\Theta_{1},..,\Theta_{M}$ as a random variable $\Theta$.
Definition 1.
For the $j$-th tree in the forest, the predicted value x will be denoted by
$m_{n,\Theta_{j},\mathcal{D}_{n}}(x)=\sum_{i=1}^{n}\frac{\mathbbm{1}_{X_{i}\in A_{n,\Theta_{j},\mathcal{D}_{n}}(x)}Y_{i}}{N_{n,\Theta_{j},\mathcal{D}_{n}}(x)}.$
Where $A_{n,\Theta_{j},\mathcal{D}_{n}}(x)$ is the cell containing $x$ and $N_{n,\Theta_{j},\mathcal{D}_{n}}(x)$ is the number of points that fall into the cell that $x$ belongs to.
For a fixed input $x\in[0,1]^{d}$, the predicted value of the tree is simply the empirical mean of the target values $Y_{i}$ for all training points $X_{i}$ that belong to the same cell as $x$. This is essentially the tree’s ”guess” for the target value corresponding to $x$, based on the training data.
Definition 2.
The finite $M$ forest is
$m_{M,n}(x)=\frac{1}{M}\sum_{j=1}^{M}m_{n,\Theta_{j},\mathcal{D}_{n}}(x).$
From a modeling point of view, we let $M\to\infty$ and consider the infinite forest estimate
$m_{\infty,n,\mathcal{D}_{n}}(x)=\mathbb{E}_{\Theta}(m_{n,\Theta,\mathcal{D}_{n}}(x)).$
The convergence holds almost surely by the law of the large numbers conditionally on the data set $\mathcal{D}_{n}.$ (Breiman) [Bre01], (Scornet) [Sco16a, Theorem 3.1].
2.1. Kernel Random Forest algorithm
In 2016, Scornet introduced kernel methods in the context of random forests, leading to the development of the kernel-based random forest algorithm known as KeRF (Kernel Random Forest) [Sco16b]. This innovative approach provided new insights and performance comparisons with traditional random forest algorithms.
To illustrate the underlying principles of KeRF, we begin by revisiting the structure of the traditional random forest algorithm. For every $x\in[0,1]^{d},$
$m_{M,n}(x)=\frac{1}{M}\sum_{j=1}^{M}\big{(}\sum_{i=1}^{n}\frac{\mathbbm{1}_{X_{i}\in A_{n,\Theta_{j},\mathcal{D}_{n}}(x)}Y_{i}}{N_{n,\Theta_{j},\mathcal{D}_{n}}(x)}\big{)}.$
The weights assigned to each observation $Y_{i}$ can be defined as
$W_{i,j,n}(x)=\frac{\mathbbm{1}_{X_{i}\in A_{n,\Theta_{j},\mathcal{D}_{n}}(x)}}{N_{n,\Theta_{j},\mathcal{D}_{n}}(x)}.$
The weight values are highly sensitive to the number of points in each cell. The influence of observations within densely populated cells is less compared to those in sparsely populated leaves. To overcome this issue, KeRF considers all tree cells containing a point simultaneously and is defined for all $x\in[0,1]^{d},$ by,
$\tilde{m}_{M,n,\Theta_{1},\Theta_{2},...,\Theta_{M}}(x)=\frac{1}{\sum_{j=1}^{M}N_{n,\Theta_{j}}(x)}\sum_{j=1}^{M}\sum_{i=1}^{n}Y_{i}\mathbbm{1}_{X_{i}\in A_{n,\Theta_{j}}(x)}.$
This way, empty cells do not affect the computation of the prediction function of the algorithm.
It is proven in [Sco16b], that this representation has indeed a kernel representation.
Proposition 1 (Scornet [Sco16b], Proposition 1).
For all $x\in[0,1]^{d}$ almost surely, it holds
$\tilde{m}_{M,n,\Theta_{1},\Theta_{2},...,\Theta_{M}}(x)=\frac{\sum_{i=1}^{n}K_{M,n}(x,X_{i})Y_{i}}{\sum_{i=1}^{n}K_{M,n}(x,X_{i})},$
where
$K_{M,n}(x,z)=\frac{1}{M}\sum_{i=1}^{M}\mathbbm{1}_{x\in A_{n,\Theta_{i},\mathcal{D}_{n}}(z)}.$
is the proximity function of the $M$ forest
The infinite random forest arises as the number of trees tends to infinity.
Definition 3.
The infinite KeRF is defined as:
$\tilde{m}_{\infty,n}(x)=\lim_{M\to\infty}\tilde{m}_{M,n}(x,\Theta_{1},\Theta_{2},...,\Theta_{M}).$
Infinite KeRF have still the kernel property,
Proposition 2 (Scornet [Sco16b], Proposition 2).
Almost surely for all $x,y\in[0.1]^{d}$
$\lim_{M\to\infty}K_{M,n}(x,y)=K_{n}(x,y),$
where
$K_{n}(x,y)=\mathbb{P}_{\Theta}(x\in A_{n}(y,\Theta)),$
is the probability that $x$ and $y$ belong to the same cell in the infinite forest.
2.2. The Centered KeRF algorithm
In this paper, we say that an estimator function $m_{n}$ of $m$ is consistent if the following $L_{2}-$type of convergence holds,
$\mathbb{E}(m_{n}(x)-m(x))^{2}\to 0,$
as $n\to\infty.$
In the centered random forest algorithm, the way the data set $\mathcal{D}_{n}$ is partitioned is independent of the data set itself.
The centered forest is designed as follows.
  • 1)
    Fix $k\in\mathbb{N}.$
  • 2)
    At each node of each individual tree choose a coordinate uniformly from $\{1,2,..d\}.$
  • 3)
    Split the node at the midpoint of the interval of the selected coordinate.
Repeat step 2)-3) $k$ times. Finally, we have $2^{k}$ leaves, or cells. A toy example of this iterative process for $k=1,2$ is in Figures 1,2. Our estimation at a point $x$ is achieved by averaging the $Y_{i}$ corresponding to the $X_{i}$ in the cell containing $x.$
Figure 1. Centered algorithm with tree level $k=1$ with the convention that $1$ corresponds to $x_{1}$ axis and $2$ to the $x_{2}$ axis.
Figure 2. Centered algorithm with tree level $k=2$ with the convention that $1$ corresponds to $x_{1}$ axis and $2$ to the $x_{2}$ axis.
Scornet in [Sco16b] introduced the corresponding kernel-based centered random forest providing explicitly the proximity kernel function.
Proposition 3.
A centered random forest kernel with $k\in\mathbb{N}$ parameter has the following multinomial expression [Sco16b, Proposition 5],
$K^{Cen}_{k}(x,z)=\sum_{\sum_{j=1}^{d}k_{j}=k}\frac{k!}{k_{1}!...k_{d}!}(\frac{1}{d})^{k}\prod_{j=1}^{d}\mathbbm{1}_{\left\lceil{2^{k_{j}}x_{j}}\right\rceil=\left\lceil{2^{k_{j}}z_{j}}\right\rceil}.$
Where $K^{Cen}_{k}$ is the Kernel of the corresponding centered random forest.
2.3. Simplified directional KeRF.
The simplified directional random forest algorithm is designed as follows:
  • 1)
    Fix $k\in\mathbb{N}.$
  • 2)
    Choose a coordinate uniformly from $\{1,2,..d\}.$
  • 3)
    For every node of each individual tree split the node at the midpoint of the interval of the preselected coordinate.
Repeat step 2)-3) $k$ times. At the end, we have $2^{k}$ leaves, or cells. A toy example of this iterative process for $k=1,2$ in Figures 3,4.
Our estimation at a point $x$ is achieved by averaging the $Y_{i}$ corresponding to the $X_{i}$ in the cell containing $x.$
For simplicity, we define the probability that two points $x,y$ are connected in the $k-$th level of a tree by $p^{sd}_{k}(x,y)$ for the simplified directional algorithm and $p_{k}(x,y)$ respectively for the centered keRF.
Figure 3. Centered algorithm with tree level $k=1$ with the convention that $1$ corresponds to $x_{1}$ axis and $2$ to the $x_{2}$ axis.
Figure 4. Centered algorithm with tree level $k=2$ with the convention that $1$ corresponds to $x_{1}$ axis and $2$ to the $x_{2}$ axis.
Theorem 1.
For every $k\in\mathbb{N}$ and every $x,y$ in $[0,1]^{d}$
$p^{sd}_{k}(x,y)=p_{k}(x,y)=K^{Cen}_{k}(x,y)$
Proof.
For $k=0,1$ the result is trivial.
We assume that for every $x,y\in[0,1]^{d}$ and for $l=0,1,...k$, $p^{sd}_{k}(x,y)=p_{k}(x,y)$ and the proof without loss of generality is provided for $d=2.$ Moreover, let $n_{sd}(k)$ resp $(n(k))$ to be the number of different tree expansions of level $k$ for the simplified directional algorithm (resp original centered algorithm), and recursively it is easy to check that
$n_{sd}(k)=2n_{sd}(k-1)=...=2^{k}$
and with the same arguments,
$n(k)=2^{k}n(k-1)=...=2^{\frac{k(k+1)}{2}}.$
Furthermore, we denote by $A_{x,y}^{k}$ the number of times that the points $x,y$ fall in the same cell in the original centered algorithm and $B_{x,y}^{k}$ for the simplified directional respectively. Then, $p^{sd}_{k}(x,y)=\frac{B_{x,y}^{k}}{n_{sd}(k)}$ and $p_{k}(x,y)=\frac{A_{x,y}^{k}}{n(k)}$ and we observe the following cases for the original centered random forest algorithm:
If $x,y$ are not connected for every possible different tree expansion of level $k$ then
$p_{k}(x,y)=p^{sd}_{k}(x,y)=p_{k+1}(x,y)=p^{sd}_{k+1}(x,y)=0.$
Furthermore, if $x,y$ are connected for some possible different tree expansion of level $k,$ but they are not connected for any tree expansions of level $k+1$ then
$p_{k+1}(x,y)=p^{sd}_{k+1}(x,y)=0.$
Next, if $x,y$ are connected for some possible different tree expansions of level $k,$ and they are connected only after a horizontal cut and not after a vertical, then
$p_{k+1}(x,y)$ $=\frac{A^{k+1}_{x,y}}{n(k+1)}$
[…]
$=\frac{1}{2}p_{k}$
$=\frac{1}{2}p^{sd}_{k}(x,y)\quad\text{by the induction hypothesis}$
and of course,
$p^{sd}_{k+1}(x,y)=\frac{B^{k+1}_{x,y}}{n_{sd}(k+1)}=\frac{1}{2}\frac{2B^{k}_{x,y}}{2^{k+1}}=\frac{1}{2}p^{sd}_{k}(x,y).$
By symmetry, the result holds as well if $x,y$ are connected for some possible different tree expansion of level $k,$ and they are connected only after a vertical cut. Finally, when $x,y$ are connected for some possible different tree expansions of level $k,$ and they are connected as well, after the next cut, in any direction then,
$p_{k+1}(x,y)$ $=\frac{A^{k+1}_{x,y}}{n(k+1)}$
$=\frac{2^{k+1}A^{k}_{x,y}}{2^{\frac{(k+1)(k+2)}{2}}}$
$=p_{k}$
$=p^{sd}_{k}(x,y)\quad\text{by the induction hypothesis}$
and,
$p^{sd}_{k+1}(x,y)=\frac{B^{k}_{x,y}}{2^{k}}=p^{sd}_{k}(x,y),$
which concludes the proof.
A simple observation from the theorem above is, that the infinite centered KeRF coincides with the infinite simplified directional KeRF. Hence, under specific assumptions, we can compute the rate of convergence of the simplified KeRF.
Corollary 1.
([IA24] Theorem 1.) Under the following assumptions:
$Y=m(X)+\epsilon$
$X\text{ is uniformly distributed on }[0,1]^{d}$
$\epsilon\sim\mathcal{N}(0,\sigma^{2}),\quad\sigma<\infty$
[…]
the rate of convergence of the simplified directional KeRF is $\mathcal{O}\left(n^{-\bigl{(}\frac{1}{1+d\log{2}}\bigr{)}}(\log{n})\right)$
3. Plots and experiments.
In this final section, we conduct numerical simulations and experiments to compare the performance of the finite-centered KeRF algorithm and the finite-simplified directional KeRF algorithm. The evaluation is carried out in terms of the $L_{2}$-error and the standard deviation of the error for various target functions $Y$. Specifically, we generated a two-dimensional feature space of size $n=1500$ comprising uniformly distributed points.
The dataset was split into training and testing subsets, with $80\%$ of the data utilized for training both algorithms, while the remaining $20\%$ was reserved for evaluation purposes computing $(\sum_{X_{i}\in\;\text{test set}}(\tilde{m}(X_{i})-Y_{i})^{2})$. To evaluate the performance of the algorithms, we considered several target functions $Y$. For each function, we trained both the finite-centered KeRF and the simplified directional KeRF on the training subset and subsequently evaluated their predictions on the remaining testing subset.
The following target functions with linear, polynomial, and exponential relationships within the feature space were considered to investigate the $L_{2}$-error with a fixed tree depth value of $k=\log_{2}{n}$ and hence, every leaf has on average $1$ data point. Moreover, the number of trees varies from $M=1,50,100,200,300,400,500$ and therefore we can empirically confirm that the two algorithms coincide asymptotically.
  1. (1)
    $Y=X_{1}+X_{2}+\epsilon$, where $\epsilon$ is a random error following a normal distribution $\mathcal{N}(0,\frac{1}{2}).$
  2. (2)
    $Y=X_{1}^{2}+X_{2}^{2}+\epsilon$, where $\epsilon$ is a random error following a normal distribution $\mathcal{N}(0,\frac{1}{2}).$
  3. (3)
    $Y=2X_{1}+e^{-X_{2}^{2}}.$
All numerical simulations were conducted using the open-source Python software https://www.python.org/,, primarily utilizing the numpy library.
(a) $L-$error for the function $Y=2X_{1}+e^{-X^{2}_{2}}$.
(b) Standard deviation for the $L-$error for the function $Y=2X_{1}+e^{-X^{2}_{2}}$.
Figure 5. Comparison of $L-$error and standard deviation.
(a) $L-$error for the function $Y=X^{2}_{1}+X^{2}_{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\frac{1}{2})$.
(b) Standard deviation for the $L-$error for the function $Y=X^{2}_{1}+X^{2}_{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\frac{1}{2})$.
Figure 6. Comparison of $L-$error and standard deviation.
(a) $L-$error for the function $Y=X_{1}+X_{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\frac{1}{2})$.
(b) Standard deviation for the $L-$error for the function $Y=X_{1}+X_{2}+\epsilon$ where $\epsilon\sim\mathcal{N}(0,\frac{1}{2})$.
Figure 7. Comparison of $L-$error and standard deviation.
As one might expect, all the figures 5(a) 6(a), 7(a) exhibit similar behavior. For small values of trees, the two algorithms demonstrate slightly different performances, however as the number of trees increases consistent with the theorem 1, both algorithms have the same performance in terms of the $L_{2}$-error consistently with 1. Similarly, the same results hold for the standard deviation of the errors 5(b) 6(b), 7(b). Overall, as it is evident from all experiments 5, 6 and 7, after $M=100$ trees the centered KeRF and the simplified directional KeRF essentially coincide.